# Duncan McCoy and Connor Sell

### v2

@@ Abstract.
We establish necessary and sufficient conditions for determining when a flat manifold can occur as a cusp cross-section within a given commensurability class of cusped arithmetic hyperbolic manifolds. This reduces the problem of identifying which commensurability classes of arithmetic hyperbolic manifolds can contain a specific flat manifold as a cusp cross-section to a question involving rational representations of the flat manifold’s holonomy group. As applications, we prove that a flat manifold $M$ with a holonomy group of odd order appears as a cusp cross-section in every commensurability class of arithmetic hyperbolic manifolds if and only if $b_{1}(M)\geq 3$. We also provide examples of flat manifolds that arise as cusp cross-sections in a unique commensurability class of arithmetic hyperbolic manifolds and exhibit examples of pairs of flat manifolds that can never appear as cusp cross-sections within the same commensurability class.
@@ 1. Introduction
The combined work of Long-Reid and McReynolds shows that every compact flat $n$-manifold arises as a cusp cross-section of hyperbolic $(n+1)$-dimensional manifold [LR02, McR09]. In particular, for a given flat manifold $B$, they construct a commensurability class of arithmetic hyperbolic manifolds such that at least one manifold in this commensurability class contains a cusp diffeomorphic to $B\times\mathbb{R}_{\geq 0}$. In general, the manifolds containing $B$ as a cusp cross-section will have many other cusps. Indeed, there are examples of flat manifolds that can not arise as the cusp cross-section of a 1-cusped hyperbolic manifold [LR00] and in sufficiently high dimensions all arithmetic manifolds necessarily have multiple cusps [Sto13]. Arithmetic hyperbolic manifolds are a particularly important class of hyperbolic manifolds because every currently known example of a hyperbolic $n$-manifold for sufficiently large $n$ is either arithmetic or is constructed directly from arithmetic pieces (for instance [GPS88]). The other known methods for constructing hyperbolic manifolds make use of hyperbolic Coxeter polytopes which exist only in dimensions $\geq 996$ [Pro86], thus limiting the dimensions in which such constructions are possible.
The principal result of this paper gives necessary and sufficient conditions for determining when a flat manifold can occur as a cusp cross-section within a given commensurability class of arithmetic hyperbolic manifolds. These conditions can be viewed as showing that every occurrence of a flat manifold as a cusp cross-section in a commensurability class of arithmetic hyperbolic manifolds is accounted for by the construction of Long-Reid and McReynolds.
Any rational quadratic form $q$ of signature $(n+1,1)$ defines an associated commensurability class of arithmetic hyperbolic $(n+1)$-manifolds. Moreover all commensurability classes of cusped arithmetic hyperbolic manifolds arise this way when $n>2$ (see Section 4). Given any compact flat $n$-manifold $B$, there is an associated finite group $H$, the holonomy group of $B$, and an associated conjugacy class of representations $\rho:H\rightarrow GL_{n}(\mathbb{Q})$, which we call the holonomy representations of $B$ (see Section 3).
@@
Suitably interpreted, the work of Long-Reid and McReynolds shows that if there is a holonomy representation of $B$ with image in $O(f;\mathbb{Z})$, where $f$ is a positive-definite $n$-dimensional rational quadratic form, then $B$ appears as a cusp cross-section in the commensurability class of hyperbolic manifolds defined by the form $f\oplus\langle 1,-1\rangle$. The main contribution of this paper is to show that if $B$ appears in the commensurability class of hyperbolic manifolds defined by a quadratic form $f\oplus\langle 1,-1\rangle$, then $B$ has a holonomy representation taking values in the orthogonal group $O(f;\mathbb{Q})$.
As a result, we obtain obstructions to flat manifolds appearing as cusp cross-sections in commensurability classes of arithmetic hyperbolic manifolds in all dimensions by studying holonomy representations. Previously, such obstructions were known only for flat 3- and 4-manifolds. The second author classified the appearance of orientable flat 3-manifolds as cusp cross-sections in commensurability classes of arithmetic hyperbolic 4-manifolds and obtained obstructions for certain flat 4-manifolds to appear as cusp cross-sections [Sel23].
@@
Since there are flat manifold with holonomy of odd order and $b_{1}(B)\leq 2$ in all dimensions, we obtain the following.
@@
Let $p$ be an odd prime and let $B$ be a flat $n$-manifold with $b_{1}(B)=0$ and holonomy group $H\cong(C_{p})^{k}$ for some $k\geq 0$. If $B$ appears in the commensurability class defined by a quadratic form $q$, then
@@
Using these manifolds, we are able to find pairs of flat manifolds that can never appear as cusp cross-sections in the same commensurability class of arithmetic hyperbolic manifolds. Consequently, we obtain restrictions on the topology of arithmetic hyperbolic manifolds.
@@
For all even $n\geq 36$ there exist pairs of flat $n$-manifolds $B_{1},B_{2}$ such that if $B_{1}$ and $B_{2}$ both appear as cusp cross-sections of a hyperbolic $(n+1)$-manifold $M$, then $M$ is not an arithmetic manifold.
@@
    $H$ is $C_{4}$ or $D_{8}$ and $\epsilon_{p}(q)=1$ for all primes $p$ such that $p\equiv 1\bmod 4$ and $-d(q)$ is a square in $\mathbb{Q}_{p}$,
@@
    $H$ is $C_{3}$, $C_{6}$, $D_{6}$ or $D_{12}$ and $\epsilon_{p}(q)=1$ for all primes $p$ such that $p\equiv 1\mod 3$ and $-d(q)$ is a square in $\mathbb{Q}_{p}$, or
@@
    $H$ is $A_{4}$, and $\epsilon_{p}(q)=1$ for all primes $p$ such that $-d(q)$ is a square in $\mathbb{Q}_{p}$.
It is interesting to note that for flat 3- and 4-manifolds the obstructions to a given flat manifold $B$ appearing in a commensurability class depend only on the isomorphism type of the holonomy group $H$. This phenomenon does not persist into higher dimensions where additional data given by the holonomy representation is needed in order to determine if a given manifold appears in a given commensurability class. For example, there are flat 5-manifolds $B$ and $B^{\prime}$ both with holonomy group $C_{3}$ such that $b_{1}(B)=1$ and $b_{1}(B^{\prime})=3$. Corollary 1.2 shows that $B^{\prime}$ appears as a cusp cross-section in every commensurability class of arithmetic manifolds, whereas $B$ does not.
@@ Acknowledgements
The first author would like to thank Steve Boyer and Michael Albanese for many helpful conversations. The second author would like to thank Alan Reid for helpful insights and for introducing him to the problem. DM is supported by an NSERC Discovery grant and a Canada Research Chair.
@@ 3. Flat manifolds
[…]
@@
[…]
@@
Since $\Gamma$ is a crystallographic group, there exists injective $\overline{\sigma}:\Gamma\rightarrow O(n)\ltimes\mathbb{R}^{n}$ with discrete image. By Lemma 3.2 this induces a holonomy representation $\sigma:H\rightarrow O(n)$ such that the diagram as in (3.2) commutes. By definition, any pair of holonomy representations are conjugate over $\mathbb{R}$ and there exists $M\in GL_{n}(\mathbb{R})$ such that $\rho=M\sigma M^{-1}$. The homomorphism $\overline{\rho}=M\overline{\sigma}M^{-1}$ provides the necessary lifting of $\rho$. ∎
@@
Let $\langle g_{1},\dots,g_{\ell}\mid r_{1},\dots,r_{m}\rangle$ be a finite presentation for $\Gamma$. Under a homomorphism $\overline{\rho}$ such that (3.4) commutes the $\rho(g_{i})$ are affine maps of the form $x\mapsto A_{i}x+\mu_{i}$ where $A_{i}=\rho(p(g_{i}))$ is determined and the $\mu_{i}\in\mathbb{R}^{n}$ satisfy the constraints imposed by the relations $r_{1},\dots,r_{m}$. Since each relation $r_{j}$ is a product of the $g_{i}$, the constraints imposed on the $\mu_{i}$ are a system of linear equations coefficients arising as the products of the $A_{i}$. In particular, this allows us to identify $\mathcal{S}$ with an algebraic subset of $\mathbb{R}^{\ell n}$ defined by linear equations over $\mathbb{Z}$. By Lemma 3.3, $\mathcal{S}$ is nonempty and contains $\overline{\rho}^{\prime}$ which is injective and has discrete image. Thus the linear equations defining $\mathcal{S}$ also admit a solution over $\mathbb{Q}$ and we may, in fact, find rational solutions arbitrarily close to $\overline{\rho}^{\prime}$. Thus there exists a homomorphism $\overline{\rho}:\Gamma\rightarrow O(f;\mathbb{Z})\ltimes\mathbb{Q}^{n}$. Furthermore, since injectivity and discreteness impose open conditions on $\mathcal{S}$, we may take $\overline{\rho}$ to be injective and with discrete image, .
@@ 4. Arithmetic Hyperbolic Manifolds
The topic of this paper is the cross-sections $B$ of the cusps of non-compact, finite-volume, arithmetic, hyperbolic $(n+1)$-manifolds. All non-compact, finite-volume, arithmetic, hyperbolic $(n+1)$-manifolds are arithmetic of simplest type over $\mathbb{Q}$ ([Mor15], Cor. 18.6.3 gives the non-trialitarian case, and an explanation for the trialitarian case based on [Tit65], [BC17], and [God63] can be found in §3.4 of [BBKS21]). For this reason, rather than giving the definition of an arithmetic group in full generality, we give a definition for those of simplest type.
@@ 4.1. Parabolic subgroups
(4.1) $\begin{split}\Phi&:O(f;\mathbb{R})\ltimes\mathbb{R}^{n}\rightarrow\Lambda_{u}\\ \Phi&(A,v)=\begin{pmatrix}A&-v&v\\ v^{T}MA&1-\frac{f(v)}{2}&\frac{f(v)}{2}\\ v^{T}MA&-\frac{f(v)}{2}&1+\frac{f(v)}{2}\end{pmatrix},\end{split}$
@@
4.2. Proof of Theorem 1.1
Theorem 1.1 will be deduced from the following statement.
@@
First, we prove that (i) implies (ii). Suppose $B$ is a cusp-cross section of an arithmetic hyperbolic manifold $M$ with associated quadratic form $q=f\oplus\langle 1,-1\rangle$. This means that there is a group $\Pi\leq O^{+}(q;\mathbb{R})$ commensurable with $O^{+}(q;\mathbb{Z})$ such that for some $y\in\partial\mathbb{H}_{q}^{n}$ we have $\Gamma=\Pi\cap\Lambda_{y}$, where $\Gamma\cong\pi_{1}(B)$ and $\Lambda_{y}$ denotes the maximal parabolic subgroup of $O^{+}(q;\mathbb{R})$ fixing $y$.
Let $T\leq\Gamma$ be the translation subgroup. First, we verify that $T\leq\Lambda_{y}\cap O^{+}(q;\mathbb{Q})$. By Lemma 4.3(ii), any $t\in T$ is a unipotent matrix. Furthermore, since $\Pi$ is commensurable with $O^{+}(q;\mathbb{Z})$, there exists an integer $k\geq 1$ such that $t^{k}\in O^{+}(q;\mathbb{Z})$. Thus Lemma 4.4 implies that $t\in O^{+}(q;\mathbb{Q})$.
Since $T\leq\Lambda_{y}\cap O^{+}(q;\mathbb{Q})$ contains a non-trivial translation, Lemma 4.3(i) implies that $y$ is a rational point of $\partial\mathbb{H}_{q}^{n}$. By Lemma 4.5 this allows us to conjugate $\Pi$ by an element of $O^{+}(q;\mathbb{Q})$ in order to assume that $\Gamma\leq\Lambda_{u}$, where $\Lambda_{u}$ is the maximal parabolic subgroup fixing $u=(0,\dots,0,1,1)^{T}$. Note that we still have $T\leq\Lambda_{u}\cap O^{+}(q;\mathbb{Q})$.
Applying $\Phi^{-1}$ from (4.1), we obtain an injective homomorphism $\Phi^{-1}:\Gamma\rightarrow O(f;\mathbb{R})\ltimes\mathbb{R}^{n}$ with discrete image, where $T$ has image in the rational translations $I\ltimes\mathbb{Q}^{n}$. Applying Lemma 3.2 to this homomorphism induces a holonomy representation of $B$ with image in $O(f;\mathbb{Q})$, as required.
Next we show that (ii) implies (iii). Suppose that $\rho$ is a holonomy representation for $B$ taking values in $O(f;\mathbb{Q})$. Since a holonomy representation is conjugate to an integral representation, then there is a matrix $T\in GL(n,\mathbb{Q})$ such that $T^{-1}\rho T$ takes values in $GL_{n}(\mathbb{Z})$. Thus $T^{-1}\rho T$ is a holonomy representation for $B$ taking values in $GL_{n}(\mathbb{Z})\cap T^{-1}O(f;\mathbb{Q})T$. However, $O(f^{\prime};\mathbb{Z})=GL_{n}(\mathbb{Z})\cap T^{-1}O(f;\mathbb{Q})T$, where $f^{\prime}$ is the form satisfying $f^{\prime}(x)=f(Tx)$ for all $x$.
Finally, we prove (iii) implies (i). This is the construction of [LR02], as interpreted in Lemma 3.4, combined with the subgroup separability results of [McR09]. If $B$ has a holonomy representation taking values in $O(f^{\prime};\mathbb{Z})$, then Lemma 3.4 combined with the isomorphism in (4.1) shows that there is a subgroup of $\Gamma\leq O^{+}(f^{\prime}\oplus\langle 1,-1\rangle;\mathbb{Z})\cap\Lambda_{u}$, where $\Gamma\cong\pi_{1}(B)$ and $\Lambda_{u}$ is the parabolic subgroup fixing the point $u=(0,\dots,0,1,1)^{T}$. McReynolds’ subgroup separability results [McR09] imply that there is a finite index torsion free subgroup of $\Pi\leq O^{+}(f^{\prime}\oplus\langle 1,-1\rangle;\mathbb{Z})$ such that $\Pi\cap\Lambda_{u}=\Gamma$. Thus $B$ arises as cusp cross-section in the commensurability class defined by $f^{\prime}\oplus\langle 1,-1\rangle$. Since $f^{\prime}\oplus\langle 1,-1\rangle$ and $f\oplus\langle 1,-1\rangle$ are rationally equivalent quadratic forms, they defined the same commensurability class of hyperbolic manifolds.
@@
Firstly Theorem 4.6 implies that if $B$ has a holonomy representation taking values in $O(f;\mathbb{Q})$ for some positive definite $f$, then $B$ appears as a cusp cross-section in the commensurability class of hyperbolic manifolds defined by $f\oplus\langle 1,-1\rangle$. Consequently $B$ appears as a cusp cross-section in the commensurability class defined by $q$ for any $q$ equivalent to $f\oplus\langle 1,-1\rangle$.
On the other hand, if $B$ appears as a cusp cross-section in the commensurability class of hyperbolic manifolds defined by $q$, then the manifolds in this class are non-compact. This implies that the form $q$ is isotropic [BHC62]. However, this implies that $q$ is equivalent to $f\oplus\langle 1,-1\rangle$ for some positive definite quadratic form $f$ of rank $n$ [Ser73, Ch. IV, Prop. 3]. Theorem 4.6 applies to show that $B$ has a holonomy representation taking values in $O(q;\mathbb{Q})$. ∎

### v3

@@ Abstract.
We establish necessary and sufficient conditions for determining when a flat manifold can occur as a cusp cross-section within a given commensurability class of cusped arithmetic hyperbolic manifolds. This reduces the problem of identifying which commensurability classes of arithmetic hyperbolic manifolds can contain a specific flat manifold as a cusp cross-section to a question involving rational representations of the flat manifold’s holonomy group. More generally we show that the holonomy representation provides an obstruction on the quasi-arithmetic manifolds containing a given flat manifold as a cusp cross-section. As applications, we prove that a flat manifold $M$ with a holonomy group of odd order appears as a cusp cross-section in every commensurability class of arithmetic hyperbolic manifolds if and only if $b_{1}(M)\geq 3$. We also provide examples of flat manifolds that arise as cusp cross-sections in a unique commensurability class of arithmetic hyperbolic manifolds and exhibit examples of pairs of flat manifolds that can never appear as cusp cross-sections in the same quasi-arithmetic hyperbolic manifold.
@@ 1. Introduction
The combined work of Long-Reid and McReynolds shows that every compact flat $n$-manifold arises as a cusp cross-section of a hyperbolic $(n+1)$-dimensional manifold [LR02, McR09]. In particular, for a given flat manifold $B$, they construct a commensurability class of arithmetic hyperbolic manifolds such that at least one manifold in this commensurability class contains a cusp diffeomorphic to $B\times\mathbb{R}_{\geq 0}$. In general, the manifolds containing $B$ as a cusp cross-section will have many other cusps. Indeed, there are examples of flat manifolds that can not arise as the cusp cross-section of a 1-cusped hyperbolic manifold [LR00] and in sufficiently high dimensions all arithmetic manifolds necessarily have multiple cusps [Sto13].
The principal result of this paper gives necessary and sufficient conditions for determining when a flat manifold can occur as a cusp cross-section within a given commensurability class of arithmetic hyperbolic manifolds. Any rational quadratic form $q$ of signature $(n+1,1)$ defines an associated commensurability class of arithmetic hyperbolic $(n+1)$-manifolds. Moreover all commensurability classes of cusped arithmetic hyperbolic manifolds arise this way when $n>2$ (see Section 4). Given any compact flat $n$-manifold $B$, there is an associated finite group $H$, the holonomy group of $B$, and an associated conjugacy class of representations $\rho:H\rightarrow GL_{n}(\mathbb{Q})$, which we call the holonomy representations of $B$ (see Section 3).
@@
Suitably interpreted, the work of Long-Reid and McReynolds shows that if there is a holonomy representation of $B$ with image in $O(f;\mathbb{Z})$, where $f$ is a positive-definite $n$-dimensional rational quadratic form, then $B$ appears as a cusp cross-section in the commensurability class of hyperbolic manifolds defined by the form $f\oplus\langle 1,-1\rangle$. The main contribution of this paper is to show that if $B$ appears in the commensurability class of hyperbolic manifolds defined by a quadratic form $f\oplus\langle 1,-1\rangle$, then $B$ has a holonomy representation taking values in the orthogonal group $O(f;\mathbb{Q})$. In fact, our work applies to the broader class of quasi-arithmetic hyperbolic manifolds.
Theorem 1.2.
Let $B$ be a compact flat $n$-manifold and let $f$ be a positive definite rational quadratic form. If $B$ appears as a cusp cross-section in a quasi-arithmetic manifold associated to the form $f\oplus\langle 1,-1\rangle$, then $B$ has a holonomy representation with image in $O(f,\mathbb{Q})$.
Since properly quasi-arithmetic manifolds exist corresponding to any rational quadratic form $q$ of rank $\geq 4$ [Tho16], Theorem 1.2 yields obstructions to the occurrence of some flat $n$-manifolds as a cusp cross-section in some non-arithmetic hyperbolic $(n+1)$-manifolds for all $n\geq 3$.
From these results, we obtain obstructions to flat manifolds appearing as cusp cross-sections in commensurability classes of arithmetic hyperbolic manifolds in all dimensions by studying holonomy representations. Previously, such obstructions were known only for flat 3- and 4-manifolds. The second author classified the appearance of orientable flat 3-manifolds as cusp cross-sections in commensurability classes of arithmetic hyperbolic 4-manifolds and obtained obstructions for certain flat 4-manifolds to appear as cusp cross-sections [Sel23].
@@
Since there are flat manifolds with holonomy of odd order and $b_{1}(B)\leq 2$ in all dimensions, we obtain the following.
@@
Let $p$ be an odd prime and let $B$ be a flat $n$-manifold with $b_{1}(B)=0$ and holonomy group $H\cong(C_{p})^{k}$ for some $k\geq 2$. If $B$ appears as a cusp cross-section in a quasi-arithmetic hyperbolic manifold corresponding to a quadratic form $q$, then
@@
Using manifolds of the form given by Theorem 1.6, we are able to obtain restrictions on the topology of quasi-arithmetic hyperbolic manifolds by showing that there are pairs of flat manifolds that can never appear as cusp cross-sections in the same quasi-arithmetic hyperbolic manifold. In particular, such cross-sections can never occur together on an arithmetic hyperbolic manifold.
@@
For all even $n\geq 36$ there exist pairs of flat $n$-manifolds $B_{1},B_{2}$ such that if $B_{1}$ and $B_{2}$ both appear as cusp cross-sections of a hyperbolic $(n+1)$-manifold $M$, then $M$ is not a quasi-arithmetic manifold.
@@
    $H$ is $C_{4}$ or $D_{8}$ and $\epsilon_{p}(q)=1$ for all primes $p$ such that $p\equiv 1\bmod 4$ and d(q)-d(q) is a square in $\mathbb{Q}_{p}$,
@@
    $H$ is $C_{3}$, $C_{6}$, $D_{6}$ or $D_{12}$ and $\epsilon_{p}(q)=1$ for all primes $p$ such that $p\equiv 1\mod 3$ and d(q)-d(q) is a square in $\mathbb{Q}_{p}$, or
@@
    $H$ is $A_{4}$, and $\epsilon_{p}(q)=1$ for all primes $p$ such that d(q)-d(q) is a square in $\mathbb{Q}_{p}$.
It is interesting to note that for flat 3- and 4-manifolds the obstructions to a given flat manifold $B$ appearing in a commensurability class depend only on the isomorphism type of the holonomy group $H$. This phenomenon does not persist into higher dimensions where the additional data of the holonomy representation is needed to determine if a flat manifold appears in a given commensurability class. For example, there are flat 5-manifolds $B$ and $B^{\prime}$ both with holonomy group $C_{3}$ such that $b_{1}(B)=1$ and $b_{1}(B^{\prime})=3$. Corollary 1.3 shows that $B^{\prime}$ appears as a cusp cross-section in every commensurability class of arithmetic manifolds, whereas $B$ does not.
@@ Acknowledgements
The first author would like to thank Steve Boyer and Michael Albanese for many helpful conversations. The second author would like to thank Alan Reid for many helpful conversations and for introducing him to the problem. DM is supported by an NSERC Discovery grant and a Canada Research Chair.
@@ 3. Flat manifolds
$1\longrightarrow\mathbb{R}^{n}\longrightarrow GL_{n}(\mathbb{R})\ltimes\mathbb{R}^{n}\overset{p}{\longrightarrow}GL_{n}(\mathbb{R})\longrightarrow 1,$
@@
$1\longrightarrow T\longrightarrow\Gamma\overset{p}{\longrightarrow}H\longrightarrow 1,$
@@
Since $\Gamma$ is a crystallographic group, there exists injective $\overline{\sigma}:\Gamma\rightarrow O(n)\ltimes\mathbb{R}^{n}$ with discrete image. By Lemma 3.2 this induces a holonomy representation $\sigma:H\rightarrow O(n)$ such that the diagram as in (3.2) commutes. By definition, any pair of holonomy representations are conjugate over $\mathbb{R}$ and there exists $M\in GL_{n}(\mathbb{R})$ such that $\rho=M\sigma M^{-1}$. The homomorphism $\overline{\rho}=M\overline{\sigma}M^{-1}$ provides the necessary lift of $\rho$. ∎
@@
Let $\langle g_{1},\dots,g_{\ell}\mid r_{1},\dots,r_{m}\rangle$ be a finite presentation for $\Gamma$. Under a homomorphism $\overline{\rho}$ such that (3.4) commutes the $\rho(g_{i})$ are affine maps of the form $x\mapsto A_{i}x+\mu_{i}$ where $A_{i}=\rho(p(g_{i}))$ is determined and the $\mu_{i}\in\mathbb{R}^{n}$ satisfy the constraints imposed by the relations $r_{1},\dots,r_{m}$. Since each relation $r_{j}$ is a product of the $g_{i}$, the constraints imposed on the $\mu_{i}$ are a system of linear equations coefficients arising as the products of the $A_{i}$. In particular, this allows us to identify $\mathcal{S}$ with an algebraic subset of $\mathbb{R}^{\ell n}$ defined by linear equations with coefficients in $\mathbb{Z}$. By Lemma 3.3, $\mathcal{S}$ is nonempty and contains $\overline{\rho}^{\prime}$ which is injective and has discrete image. Thus the linear equations defining $\mathcal{S}$ also admit a solution over $\mathbb{Q}$ and we may, in fact, find rational solutions arbitrarily close to $\overline{\rho}^{\prime}$. Thus there exists a homomorphism $\overline{\rho}:\Gamma\rightarrow O(f;\mathbb{Z})\ltimes\mathbb{Q}^{n}$. Furthermore, since injectivity and discreteness impose open conditions on $\mathcal{S}$, we may take $\overline{\rho}$ to be injective and with discrete image.
@@ 4. Arithmetic hyperbolic manifolds
The topic of this paper is the cross-sections $B$ of the cusps of non-compact, finite-volume, arithmetic, hyperbolic $(n+1)$-manifolds. All non-compact, finite-volume, arithmetic, hyperbolic $(n+1)$-manifolds are arithmetic of simplest type over $\mathbb{Q}$ ([Mor15], Cor. 18.6.3 gives the non-trialitarian case, and an explanation for the trialitarian case based on [Tit66], [BC17], and [God64] can be found in §3.4 of [BBKS21]). For this reason, rather than giving the definition of an arithmetic group in full generality, we give a definition for those of simplest type. For similar reasons, we will also only define quasi-arithmetic manifolds analogous to arithmetic ones of simplest type.
@@
For the purposes of this article, the following definition of quasi-arithmetic hyperbolic manifold will be sufficient. A more general definition can be found in [Vin67].
Definition 4.3.
Let $q$ be a rational quadratic form of signature $(n+1,1)$. A quasi-arithmetic hyperbolic manifold associated to $q$ is a finite-volume hyperbolic manifold of the form $\mathbb{H}_{q}^{n+1}/\Gamma$ where $\Gamma\leq O^{+}(q;\mathbb{R})$ is discrete, torsion-free and commensurable to a subgroup of $O^{+}(q;\mathbb{Q})$.
@@ 4.1. Parabolic subgroups
[…]
@@
4.2. Proof of Theorem 1.1 and Theorem 1.2
First, we show that the existence of a cusp cross-section implies the existence of a holonomy representation.
Theorem 1.2.
Let $B$ be a compact flat $n$-manifold and let $f$ be a positive definite rational quadratic form. If $B$ appears as a cusp cross-section in a quasi-arithmetic manifold associated to the form $f\oplus\langle 1,-1\rangle$, then $B$ has a holonomy representation with image in $O(f,\mathbb{Q})$.
Proof.
If $B$ appears as a cusp cross-section in such a quasi-arithmetic manifold, then there is a group $\Gamma\leq O^{+}(q;\mathbb{R})$ commensurable with a $O^{+}(q;\mathbb{Q})$ such that for some $y\in\partial\mathbb{H}_{q}^{n}$ we have $\Pi=\Gamma\cap\Lambda_{y}$, where $\Pi\cong\pi_{1}(B)$ and $\Lambda_{y}$ denotes the maximal parabolic subgroup of $O^{+}(q;\mathbb{R})$ fixing $y$.
Let $T\leq\Pi$ be the translation subgroup. First, we verify that $T\leq\Lambda_{y}\cap O^{+}(q;\mathbb{Q})$. By Lemma 4.4(ii), any $t\in T$ is a unipotent matrix. Furthermore, since $\Gamma$ is commensurable with a subgroup of $O^{+}(q;\mathbb{Q})$, there exists an integer $k\geq 1$ such that $t^{k}\in O^{+}(q;\mathbb{Q})$. Thus Lemma 4.5 implies that $t\in O^{+}(q;\mathbb{Q})$.
Since $T\leq\Lambda_{y}\cap O^{+}(q;\mathbb{Q})$ contains a non-trivial translation, Lemma 4.4(i) implies that $y$ is a rational point of $\partial\mathbb{H}_{q}^{n}$. By Lemma 4.6 this allows us to conjugate $\Gamma$ by an element of $O^{+}(q;\mathbb{Q})$ in order to assume that $\Pi\leq\Lambda_{u}$, where $\Lambda_{u}$ is the maximal parabolic subgroup fixing $u=(0,\dots,0,1,1)^{T}$. Note that we still have $T\leq\Lambda_{u}\cap O^{+}(q;\mathbb{Q})$.
Applying $\Phi^{-1}$ from (4.1), we obtain an injective homomorphism $\Phi^{-1}:\Pi\rightarrow O(f;\mathbb{R})\ltimes\mathbb{R}^{n}$ with discrete image, where $T$ has image in the rational translations $I\ltimes\mathbb{Q}^{n}$. Applying Lemma 3.2 to this homomorphism induces a holonomy representation of $B$ with image in $O(f;\mathbb{Q})$, as required. ∎
We make note of the following more elaborate statement from which Theorem 1.1 is easily deduced.
@@
[…]
@@
Firstly, Theorem 4.7 implies that if $B$ has a holonomy representation taking values in $O(f;\mathbb{Q})$ for some positive definite $f$, then $B$ appears as a cusp cross-section in the commensurability class of hyperbolic manifolds defined by $f\oplus\langle 1,-1\rangle$. Consequently $B$ appears as a cusp cross-section in the commensurability class defined by $q$ for any $q$ equivalent to $f\oplus\langle 1,-1\rangle$.
On the other hand, if $B$ appears as a cusp cross-section in the commensurability class of hyperbolic manifolds defined by $q$, then the manifolds in this class are non-compact. This implies that the form $q$ is isotropic [BHC62]. However, this implies that $q$ is equivalent to $f\oplus\langle 1,-1\rangle$ for some positive definite quadratic form $f$ of rank $n$ [Ser73, Ch. IV, Prop. 3]. Theorem 4.7 applies to show that $B$ has a holonomy representation taking values in $O(f;\mathbb{Q})$. ∎